# A Differentiable Alignment Framework for Sequence-to-Sequence Modeling via Optimal Transport

### v1

@@ 4.2.1 Sequences-to-Sequences Distance
[…]
where $\bm{\alpha}^{*}$ is a minimum for which $\mathcal{S}_{r}(\{{\bm{x}}\}_{n},\{{\bm{y}}\}_{m})=0$. It should be noted that this condition holds also when $C$ is neither symmetric nor satisfies the triangular inequality, but is separated (like the cross-entropy for example).
Proof. See Appendix A.2.5.
The consequence of the previous proposition is that we can learn a transformation through gradient descent using a trainable network $F$ which maps input sequences $\{{\bm{x}}\}_{n}$ to target sequences $\{{\bm{y}}\}_{m}$ (with $n\geq m$) by solving the optimization problem:
$\min_{F}\mathcal{S}_{r}(F(\{{\bm{x}}\}_{n}),\{{\bm{y}}\}_{m})$ (8)
@@
Note on Dynamic Time Warping (DTW): It is important to highlight the distinction between our approach and DTW-based (Itakura, 1975) alignment methods, particularly the differentiable variations such as soft-DTW (Cuturi & Blondel, 2018). These methods generally have quadratic complexity (Cuturi & Blondel, 2018), making them significantly more computationally expensive than ours. Furthermore, in DTW-based methods, the alignment emerges as a consequence of the sequences themselves. When the function $F$ is powerful, the model can collapse by generating a sequence $F(\{{\bm{x}}\}_{n})$ that induces a trivial alignment (Haresh et al., 2021) (see Appendix A.3.1, where we conducted experiments using soft-DTW for ASR to illustrate this). To mitigate this issue, regularization losses (Haresh et al., 2021; Meghanani & Hain, 2024) or constraints on the capacity of $F$ (Vayer et al., 2022; Zhou & la Torre, 2009) are commonly introduced. However, using regularization losses lacks theoretical guarantees and introduces additional hyperparameters. Furthermore, constraining the capacity of $F$, although more theoretically sound, makes tasks requiring powerful encoders on large datasets impractical. In contrast, our method decouples the computation of the alignment from the transformation function $F$, offering more flexibility to the model as well as built-in temporal alignment constraints and theoretical guarantees against collapse.
@@ 4.3 Application to ASR: OTTC Loss
$\small F(\{{\bm{x}}\}_{n})=\{[p_{l_{1}}({\bm{x}}_{i}),\dots,p_{l_{|L|+1}}({\bm{x}}_{i})]^{T}\}_{i=1}^{n}.$ (9)
The alignment between $F(\{{\bm{x}}\}_{n})$ and $\{{\bm{y}}\}_{m}$ is parameterized by $\bm{\alpha}[\{{\bm{x}}\}\textdegree_{n},W]\in\Delta^{n}$, defined as
@@ 4.4 Link with CTC Loss
In this section, we link the CTC and the proposed OTTC losses. In the context of CTC, we denote by $\mathcal{B}$ the mapping which reduces any sequences by deleting repeated vocabulary (similarly to the previously defined $\mathcal{A}$ mapping in Proposition 5) and then deleting the blank token $\phi$ (e.g., ($\mathcal{B}(\{GGOO\phi ODD\})=\{GOOD\}$). The objective of CTC is to maximise the probability of all possible paths $\{\bm{\pi}\}_{n}$ of length $n$ through minimizing
$-\mkern-25.0mu\sum_{\{\bm{\pi}\}_{n}\in\mathcal{B}^{-1}(\{{\bm{y}}\}_{m})}\log p(\{\bm{\pi}\}_{n})=-\mkern-25.0mu\sum_{\{\bm{\pi}\}_{n}\in\mathcal{B}^{-1}(\{{\bm{y}}\}_{m})}\log\prod_{i=1}^{n}p(\bm{\pi}_{i}),$ (12)
@@
Let us consider a path $\{\bm{\pi}\}_{n}\in\mathcal{B}^{-1}(\{{\bm{y}}\}_{m})$. Such a path can be seen as an alignment (see Figure 3), where $\{{\bm{x}}_{i}\}$ and $\{{\bm{y}}_{j}\}$ are aligned iff $\bm{\pi}_{i}={\bm{y}}_{j}$. By denoting $\bm{A}_{\pi}$ as the corresponding discrete monotonic alignment, one can write:
$\log p(\{\bm{\pi}\}_{n})=\sum_{i=1}^{n}\log p_{\bm{\pi}_{i}}({\bm{x}}_{i})$ $=\>\>\>\>\>-\mkern-20.0mu\sum_{\begin{subarray}{c}i,j=1\\ (i,j)\in\mathbf{A}_{\bm{\pi}}\end{subarray}}^{n,m}\mkern-10.0muC_{e}(\bm{\pi}_{j},{\bm{y}}_{i})$
=$\mkern-20.0mu\stackrel{{\scriptstyle\exists\bm{\alpha}\in\Delta^{n}}}{{=}}-\mkern-40.0mu\sum_{\begin{subarray}{c}i,j=1\\ \bm{\gamma}_{p}^{n,\bm{\beta}_{m}}(\bm{\alpha})_{i,j}>0\end{subarray}}^{n,m}\mkern-35.0muC_{e}(\bm{\pi}_{j},{\bm{y}}_{i}),$ (13)
@@
The continuous relaxation (i.e. making the problem continuous with respect to alignment) of the last term in this sequence of equalities results in $-\mathcal{L}_{OTTC}$. Therefore, OTTC can be seen as relaxation of the probability associated with a single path, enabling a differentiable path search mechanism. Essentially, OTTC optimization focuses on maximizing the probability of exactly one path, in contrast to CTC, which maximizes the probability across all valid paths.
Additionally, OTTC does not incentivize paths containing many blank tokens, unlike CTC. In CTC, the peaky behavior arises because maximizing the marginal probability over all valid paths can incentivize the model to assign more frames to the blank token (Zeyer et al., 2021). In contrast, OTTC does not rely on a blank token to indicate that a frame $i$ should not be classified (blank tokens are only used to separate consecutive tokens). Instead, the model simply sets the corresponding weight $\alpha_{i}$ to 0 (see Figure 2). This mechanism avoids the peaky behavior exhibited by CTC.
@@ 5 Experimental Setup
To demonstrate the viability of the proposed OTTC loss framework, we conduct several proof-of-concept experiments on the ASR task. To this end, we compare alignment quality and ASR performance using the proposed OTTC framework and the existing CTC-based model. Note that an efficient batched implementation of OTTC along with the full code to reproduce our experimental results will be made publicly available.
Datasets.  We conduct our experiments on popular open-source datasets, i.e., the TIMIT (Garofolo et al., 1993), AMI (Carletta et al., 2005), and LibriSpeech (Panayotov et al., 2015). TIMIT is a 5-hour English dataset with time-aligned transcriptions, including exact time-frame phoneme transcriptions, making it a standard benchmark for ASR and phoneme segmentation tasks. We report results on the standard eval set. AMI is an English spontaneous meeting speech corpus that serves as a good benchmark to evaluate our approach in a realistic conversational scenario, due to its spontaneous nature and prior use in alignment evaluation (Rastorgueva et al., 2023). For our experiments on this dataset, we train models on the individual head microphone (IHM) split comprising 80 hours of audio, and report results on the official eval set. LibriSpeech is an English read-speech corpus derived from audiobooks, containing 1000 hours of data. It is a standard benchmark for reporting ASR results. For our experiments, we train models on the official 100-hour, 360-hour, and 960-hour splits, and report results on the two official test sets.
Table 1: Alignment performance of the CTC loss-based ASR model and our proposed OTTC loss-based ASR model on the TIMIT and AMI datasets. Peaky behavior is measured as the percentage of frames assigned to blank or space symbols (lower is better) - For TIMIT, we subtract the percentage of real silence, as it is available, unlike in AMI. The F1 score reflects the accuracy of the starting frame of the predicted tokens. IDR quantifies the overlap between reference and predicted word segments (higher is better).
@@
Peaky F1 Score IDR Peaky F1 Score IDR
@@
Model architecture.  We use the 300M parameter version of the well-known Wav2Vec2-large (Baevski et al., 2020) as the base model for acoustic embeddings in all the experiments conducted in this work. The Wav2Vec2 is a self-supervised model pre-trained on 60K hours of unlabeled English speech. For the baseline CTC-based models, we stack a dropout layer followed by a linear layer for logits prediction, termed the logits prediction head. For the proposed OTTC loss based models, we use a dropout and a linear layer (identical to the baseline) for logits prediction. In addition, as described in Section 4.3, we apply a dropout layer followed by two linear layers on top of the Wav2Vec2-large model for OT weight prediction, with a GeLU (Hendrycks & Gimpel, 2016) non-linearity in between, termed the OT weights prediction head. Note that the output from the Wav2Vec2-large model is used as input for both the logit and OT weight prediction heads, and the entire model is trained using the OTTC loss.
@@
Performance metrics.  ASR performance is evaluated using the standard WER. Alignment quality is assessed using three metrics, i.e., peaky behavior, starting frame accuracy, and Intersection Duration Ratio (IDR). Peaky behavior in CTC models is characterized by a large proportion of audio frames being assigned to blank or space symbols (non-alphabet symbols) (Zeyer et al., 2021). To quantify this, we compute the average percentage of frames mapped to these symbols. Starting frame accuracy is evaluated using the F1 score, following the methodology in (Rastorgueva et al., 2023). On TIMIT, where ground truth alignments are available, we assess timestamps at the phoneme level. In contrast, on AMI, where ground truth alignments are unavailable, we use forced alignment as in (Rastorgueva et al., 2023), but restrict our evaluation to word-level timestamps which are generally more reliable than phoneme-, letter-, or subword-level timestamps. Please note that the F1 score used in the state-of-the-art literature is based only on the starting frame of the predicted token, which does not fully reflect alignment quality. To address this, we introduce IDR, which measures the overlap between predicted and reference word segments, normalized by the reference duration. This provides a finer-grained assessment of temporal alignment.
@@
Training details.  In all our experiments, we use the AdamW optimizer (Loshchilov & Hutter, 2019) for training. For TIMIT and LibriSpeech, the initial learning rate is set to $lr\!=2e^{-4}$, with a linear warm-up for the first $500$ steps followed by a linear decay until the end of training. For AMI, the initial learning rate is set to $lr\!=1.25e^{-3}$, with a linear warm-up during the first $10\%$ of the steps, also followed by linear decay. We train both CTC-based and OTTC-based models for 40 epochs, reporting the test set WER at the final epoch. In our OTTC-based models, both the logits and OT weight prediction heads are trained for the first 30 epochs. During the final 10 epochs, the OT weight prediction head is fixed, while training continues on the logits prediction head. For experiments on the LibriSpeech (resp. TIMIT) dataset, we use character-level (resp. phoneme-level) tokens to encode text. Given the popularity of subword-based units for encoding text (Sennrich et al., 2016), we sought to observe the behavior of OTTC-based models when tokens are subword-based, where a token can contain more than one character. For the experiments on the AMI dataset, we use the SentencePiece tokenizer (Kudo & Richardson, 2018) to train subwords from the training text. Greedy decoding is used for both the CTC and OTTC models to generate the hypothesis text.
@@ 6 Results and Discussion
Alignment quality.  We begin by analyzing the alignment performance of the models on the TIMIT and AMI datasets, with results shown in Table 1. Our proposed OTTC model consistently outperforms the CTC-based models across all alignment metrics on both the datasets. A key observation is the significant difference in the percentage of frames assigned to non-alphabet symbols by the CTC models, which highlights the peaky behavior inherent in these models. Specifically, CTC models tend to assign a large proportion of frames to the blank or space symbols, reflecting a misalignment in predicting word boundaries. In contrast, the OTTC model avoids this issue, preventing the extreme peaky behavior observed in CTC models. While the OTTC model also outperforms the CTC model in F1 score, the margin of improvement is smaller. However, the IDR reveals a substantial advantage for OTTC, with a significant improvement over CTC. This indicates that CTC models often either delay the prediction of word starts or assigns too few frames to non-blank symbols, reinforcing the peaky behavior. Additionally, the performance improvement on the AMI dataset is particularly significant, given its nature of spontaneous meeting speech. This demonstrates how effectively the OTTC loss adapts to varying speaking rates, showcasing the robustness of our framework in learning alignments despite speech variability.
Table 2: Word Error Rate (WER%) comparison between the CTC loss-based ASR model and our proposed OTTC loss-based ASR model on the TIMIT, AMI, and LibriSpeech datasets. Lower WER is better.
@@
WER.  Next, we report the ASR performance in terms of WER in Table 2. On the TIMIT dataset, the OTTC model shows a slightly higher WER compared to the CTC model, and while the performance gap is larger on the AMI dataset, it’s encouraging to observe consistent performance despite the varied nature of speech. On the LibriSpeech dataset, using the 100-hour training split, the OTTC model achieves a WER of $3.77\%$ on test-clean. As we scale the training dataset (100h $\rightarrow$ 360h $\rightarrow$ 960h), we observe a monotonic improvement in WER for the proposed OTTC-based models, similarly to the CTC-based models. Although the WERs achieved by the OTTC-based models are typically higher than the CTC-based models, the presented results underscore the experimental validity of the SOTD as a metric and demonstrate that learning a single alignment can yield promising results in E2E ASR.
@@
Qualitative alignment comparison.  Apart from quantitative alignment comparison (Table 1), we show an alignment from the CTC- and OTTC-based models in Figure 4. For CTC, it can be seen that the best path aligns most frames to the blank token, resulting in a peaky behavior (Zeyer et al., 2021). In contrast, the OTTC model learns to align all frames to non-blank tokens. This effectively mitigates the peaky behavior observed in the CTC model. Note that OTTC allows dropping frames during alignment (see Section 4.4), however, in practice, we observed that only a few frames are dropped. For additional insights, we plot the evolution of the alignment for the OTTC model during the course of training in Figures 5 & 7. It is evident that the alignment learned early in the training process remains relatively stable as training progresses. The most notable changes occur at the extremities of the predicted label clusters. This observation led us to the decision to freeze the OT weight predictions for the final 10 epochs, otherwise, even subtle changes in alignment could adversely impact the logits predictions because same base model is shared for predicting both the logits and the alignment OT weights.

### v2

@@ 4.2.1 Sequence-to-Sequence Distance
$\mathcal{S}_{r}(\{{\bm{x}}\}_{n},\{{\bm{y}}\}_{m})=0\text{\>\>iff\>\>}\mathcal{A}(\mathcal{P}_{\alpha^{*}}(\{{\bm{x}}\}_{n}))=\mathcal{A}(\{{\bm{y}}\}_{m}),$ (7)
where $\bm{\alpha}^{*}$ is a minimum for which $\mathcal{S}_{r}(\{{\bm{x}}\}_{n},\{{\bm{y}}\}_{m})=0$. It should be noted that this condition holds also when $C$ is neither symmetric nor satisfies the triangular inequality, but is separated (like the cross-entropy for example). (Proof. See Appendix A.2.5. )
The consequence of the previous proposition is that we can learn a transformation through gradient descent using a trainable network $F$ which maps input sequences $\{{\bm{x}}\}_{n}$ to target sequences $\{{\bm{y}}\}_{m}$ (with $n\geq m$) by solving the optimization problem
$\min_{F}\mathcal{S}_{r}(F(\{{\bm{x}}\}_{n}),\{{\bm{y}}\}_{m}).$ (8)
@@
Note on Dynamic Time Warping (DTW): A note on the distinction between our approach and DTW-based methods Itakura1975MinimumPR can be found in Appendix A.4.
@@ 4.3 Application to ASR: OTTC Loss
$F(\{{\bm{x}}\}_{n})=\{[p_{l_{1}}({\bm{x}}_{i}),\dots,p_{l_{|L|+1}}({\bm{x}}_{i})]^{T}\}_{i=1}^{n}.$ (9)
The alignment between $F(\{{\bm{x}}\}_{n})$ and $\{{\bm{y}}\}_{m}$ is parameterized by $\bm{\alpha}[\{{\bm{x}}\}_{n},W]\in\Delta^{n}$, defined as
@@ 4.4 Link with CTC Loss
In this section, we link the CTC and the proposed OTTC losses. In the context of CTC, we denote by $\mathcal{B}$ the mapping which reduces any sequences by deleting repeated vocabulary (similarly to the previously defined $\mathcal{A}$ mapping in Proposition 5) and then deleting the blank token $\phi$ (e.g., $\mathcal{B}(\{GGOO\phi ODD\})=\{GOOD\}$). The objective of CTC is to maximise the probability of all possible paths $\{\bm{\pi}\}_{n}$ of length $n$ through minimizing
$-\log\mkern-25.0mu\sum_{\{\bm{\pi}\}_{n}\in\mathcal{B}^{-1}(\{{\bm{y}}\}_{m})}p(\{\bm{\pi}\}_{n})=-\log\mkern-25.0mu\sum_{\{\bm{\pi}\}_{n}\in\mathcal{B}^{-1}(\{{\bm{y}}\}_{m})}\prod_{i=1}^{n}p(\bm{\pi}_{i}),$ (12)
@@
Let us consider a path $\{\bm{\pi}\}_{n}\in\mathcal{B}^{-1}(\{{\bm{y}}\}_{m})$. Such a path can be seen as an alignment (see Figure 3), where $\{{\bm{x}}_{i}\}$ and $\{{\bm{y}}_{j}\}$ are aligned iff $\bm{\pi}_{i}={\bm{y}}_{j}$. By denoting $\bm{A}_{\pi}$ as the corresponding discrete monotonic alignment, one can write
$-\log p(\{\bm{\pi}\}_{n})=-\sum_{i=1}^{n}\log p_{\bm{\pi}_{i}}({\bm{x}}_{i})$ $=\>\>\>\>\>\sum_{\begin{subarray}{c}i,j=1\\ (i,j)\in\mathbf{A}_{\bm{\pi}}\end{subarray}}^{n,m}C_{e}(\bm{\pi}_{j},{\bm{y}}_{i})$ $\mkern-20.0mu\stackrel{{\scriptstyle\exists\bm{\alpha}\in\Delta^{n}}}{{=}}\sum_{\begin{subarray}{c}i,j=1\\ \bm{\gamma}_{p}^{n,\bm{\beta}_{m}}(\bm{\alpha})_{i,j}>0\end{subarray}}^{n,m}C_{e}(\bm{\pi}_{j},{\bm{y}}_{i}).$ (13)
@@
The continuous relaxation (i.e., making the problem continuous with respect to alignment) of the last term in this sequence of equalities results in $\mathcal{L}_{OTTC}$. Therefore, OTTC can be seen as relaxation of the probability associated with a single path, enabling a differentiable path search mechanism. Essentially, OTTC optimization focuses on maximizing the probability of exactly one path, in contrast to CTC, which maximizes the probability across all valid paths.
Additionally, OTTC does not incentivize paths containing many blank tokens, unlike CTC. In CTC, the peaky behavior arises because maximizing the marginal probability over all valid paths can incentivize the model to assign more frames to the blank token ctc-peaky . In contrast, OTTC does not rely on a blank token to indicate that a frame $i$ should not be classified (blank tokens are only used to separate consecutive tokens). Instead, the model simply sets the corresponding weight $\alpha_{i}$ to 0 (see Figure 2). This mechanism avoids the peaky behavior exhibited by CTC.
@@ 5 Experimental Setup
To demonstrate the viability of the proposed OTTC loss framework, we conduct several proof-of-concept experiments on the ASR task. To this end, we compare alignment quality and ASR performance using the proposed OTTC framework and existing CTC-based models. Note that an efficient batched implementation of OTTC along with the full code to reproduce our experimental results will be made publicly available.
Datasets.  We conduct our experiments on popular open-source datasets, i.e., the TIMIT garofolo1993timit , AMI ami , and LibriSpeech panayotov2015librispeech . TIMIT is a 5-hour English dataset with time-aligned transcriptions, including exact time-frame phoneme transcriptions, making it a standard benchmark for ASR and phoneme segmentation tasks. We report results on the standard eval set. AMI is an English spontaneous meeting speech corpus that serves as a good benchmark to evaluate our approach in a realistic conversational scenario, due to its spontaneous nature and prior use in alignment evaluation rastorgueva23_interspeech . For our experiments on this dataset, we train models on the individual head microphone (IHM) split comprising 80 hours of audio, and report results on the official eval set. LibriSpeech is an English read-speech corpus derived from audiobooks, containing 1000 hours of data. It is a standard benchmark for reporting ASR results. For our experiments, we train models on the official 100-hour, 360-hour, and 960-hour splits, and report results on the two official test sets.
Baselines.  We benchmark our performance against the standard CTC. To specifically compare alignment quality, particularly regarding the mitigation of the peaky behavior inherent in CTC-based models, we also include CR-CTC yao2024cr . CR-CTC serves as a strong baseline, chosen for its established effectiveness against such peaky alignments.
Table 1: Alignment performance of the CTC-, CR-CTC-, and OTTC-based ASR models on the TIMIT and AMI datasets. For TIMIT, we subtract the percentage of real silence, as it is available, unlike in AMI.
@@
Peaky($\downarrow$) F1 Score ($\uparrow$) IDR ($\uparrow$) Peaky ($\downarrow$) F1 Score ($\uparrow$) IDR ($\uparrow$)
@@
Model architectures.  We use the 300M parameter version of the well-known Wav2Vec2-large baevski2020wav2vec as the base model for acoustic embeddings in all the experiments conducted in this work. The Wav2Vec2 is a self-supervised model pre-trained on 60K hours of unlabeled English speech. For the baseline CTC-based models, we stack a dropout layer followed by a linear layer for logits prediction, termed the logits prediction head. For the proposed OTTC loss based model, we use a dropout and a linear layer (identical to the baseline) for logits prediction. In addition, as described in Section 4.3, we apply a dropout layer followed by two linear layers on top of the Wav2Vec2-large model for OT weight prediction, with a GeLU hendrycks2016gaussian non-linearity in between, termed the OT weights prediction head. Note that the output from the Wav2Vec2-large model is used as input for both the logit and OT weight prediction heads, and the entire model is trained using the OTTC loss.
@@
Performance metrics.  Alignment quality is assessed using three metrics: peaky behavior, starting frame accuracy, and Intersection Duration Ratio (IDR). Peaky behavior, a common characteristic of CTC-based models, refers to a large proportion of audio frames being assigned to blank or space symbols (non-alphabet symbols) ctc-peaky . To quantify this, we compute the average percentage of frames mapped to these symbols. Starting frame accuracy is evaluated using the F1 score, following the methodology proposed in rastorgueva23_interspeech . It is important to note that this F1 score reflects only the correctness of the predicted token’s starting frame and does not fully capture alignment quality. To address this, we introduce IDR, which measures the overlap between predicted and reference word segments, normalized by the reference duration. This provides a finer-grained assessment of temporal alignment. These alignment metrics are computed only on the TIMIT and AMI datasets due to the lack of reliable ground-truth or forced-alignment annotations for LibriSpeech. On TIMIT, where ground-truth alignments are available, we assess alignment at the phoneme level. For AMI, which lacks ground-truth timestamps, we follow the forced-alignment approach in rastorgueva23_interspeech , but restrict evaluation to word-level timestamps, as they are generally more reliable than phoneme-, letter-, or subword-level annotations. Finally, ASR performance is evaluated using the WER on all considered databases.
@@
Training details.  In all our experiments, we use the AdamW optimizer loshchilov2018decoupled for training. For TIMIT and LibriSpeech, the initial learning rate is set to $lr\!=2e^{-4}$, with a linear warm-up for the first $500$ steps followed by a linear decay until the end of training. For AMI, the initial learning rate is set to $lr\!=1.25e^{-3}$, with a linear warm-up during the first $10\%$ of the steps, also followed by linear decay. We train all considered models for 40 epochs, reporting the test set WER at the final epoch. In our OTTC-based models, both the logits and OT weight prediction heads are trained for the first 30 epochs. During the final 10 epochs, the OT weight prediction head is fixed, while training continues on the logits prediction head. For experiments on the LibriSpeech (resp. TIMIT) dataset, we use character-level (resp. phoneme-level) tokens to encode text. Given the popularity of subword-based units for encoding text sennrich2016neural , we sought to observe the behavior of OTTC-based models when tokens are subword-based, where a token can contain more than one character. For the experiments on the AMI dataset, we use the SentencePiece tokenizer kudo2018sentencepiece to train subwords from the training text. Greedy decoding is used for all considered models to generate the hypothesis text.
@@ 6 Results and Discussion
Alignment quality.  We begin by analyzing the alignment performance of the models on the TIMIT and AMI datasets, with results shown in Table 1. Our proposed OTTC model consistently outperforms the CTC-based models across all alignment metrics on both datasets. A key observation is the significant difference in the percentage of frames assigned to non-alphabet symbols by the CTC-based models, highlighting the peaky behavior inherent in these models. Specifically, the baseline CTC-based models tend to assign a large proportion of frames to blank or space symbols, reflecting a misalignment in predicted word boundaries. In contrast, the OTTC model avoids this issue, preventing extreme peaky behavior observed in CTC-based models. While the OTTC model also outperforms the CTC-based models in F1 score, the margin of improvement is smaller. However, the IDR reveals a substantial advantage for OTTC, with a significant improvement over CTC and CR-CTC. This indicates that CTC-based models often either delay the prediction of word starts or assigns too few frames to non-blank symbols, reinforcing the peaky behavior. Additionally, the performance improvement on the AMI dataset is particularly significant, given its nature of meeting speech. This demonstrates how effectively the OTTC loss adapts to varying speaking rates, showcasing the robustness of our framework in learning alignments despite speech variability.
Table 2: Word Error Rate (WER%) comparison between the baseline CTC model and the proposed OTTC model on all considered datasets. Lower WER is better.
@@
WER.  ASR performance in terms of WER for the CTC model and the proposed OTTC model is depicted in Table 2 for all considered datasets. On the TIMIT dataset, the OTTC model shows a slightly higher WER compared to the CTC model, and while the performance gap is larger on the AMI dataset, it’s encouraging to observe consistent performance despite the varied nature of speech. On the LibriSpeech dataset, using the 100-hour training split, the OTTC model achieves a WER of $3.77\%$ on test-clean. As we scale the training dataset (100h $\rightarrow$ 360h $\rightarrow$ 960h), we observe a monotonic improvement in WER for the proposed OTTC-based models, similarly to the CTC-based models. Although the WERs achieved by the OTTC-based models are typically higher than the CTC-based models, the presented results underscore the experimental validity of the SOTD as a metric and demonstrate that learning a single alignment can yield promising results in E2E ASR.
Qualitative alignment comparison.  Apart from quantitative alignment comparison (Table 1), we show an alignment from the CTC- and OTTC-based models in Figure 4.
@@
For CTC, it can be seen that the best path aligns most frames to the blank token, resulting in peaky behavior ctc-peaky . In contrast, the OTTC model learns to align all frames to non-blank tokens. This effectively mitigates the peaky behavior observed in the CTC model. Note that OTTC allows dropping frames during alignment (see Section 4.4), however, in practice, we observed that only a few frames are dropped. For additional insights, we plot the evolution of the alignment for the OTTC model during the course of training in Figures 6 & 7. It is evident that the alignment learned early in the training process remains relatively stable as training progresses. The most notable changes occur at the extremities of the predicted label clusters. This observation led us to the decision to freeze the OT weight predictions for the final 10 epochs, otherwise, even subtle changes in alignment could adversely impact the logits predictions because same base model is shared for predicting both the logits and the alignment OT weights.